# Low-rank matrix and tensor approximations: advancing efficiency of machine-learning interatomic potentials

### v1

@@ III.2 Determining the optimal ranks on the example of Mo-Nb-Ta-W
Figure 5: Loss function and root mean square errors (RMSEs) for energies and forces for different MTPs trained on Mo-Nb-Ta-W with the optimal ranks. The calculations are conducted on the validation set. We provide the results with 68% confidence interval (i.e., 1-$\sigma$ interval).
Figure 6: Loss function and root mean square errors (RMSEs) for energies and forces for different potentials trained on Mo-Nb-Ta-W with 50% compression. The calculations are conducted on the validation set. We provide the results with 68% confidence interval (i.e., 1-$\sigma$ interval).
III.3 Loss functions and training errors
@@
In this subsection, we analyze the values of loss functions and fitting errors calculated with ensembles of different MTPs with the optimal ranks and with the compression of approximately 50% (i.e., two times reduction) of the number of the radial parameters in the base MTP.
@@
The results of fitting of potentials on Mo-Nb-Ta-W are shown in Figures  5 and  6. In Figure  5 we observe the decrease in loss functions, energy and force fitting errors with the increase of the level of all MTPs. We also see that all the compressed MTPs (MF, MFRA, R-MF, and TF) of the 16th and 18th levels give higher accuracy, i.e., smaller loss functions, energy and force fitting errors than the base MTPs. However, for the MTPs of the 12th level, we conclude that the compressed MTPs give either worse (R-MF) or slightly lower (MF, TF, and MFRA) results compared to the base MTP. This effect can be related to the number of linear parameters and the complexity of MTP. There are 29 linear parameters and three radial functions in the MTP of the 12th level, and thus, all the radial parameters presented in the base MTP can be critical for accurate fitting of the compressed MTPs. However, MTPs of the 16th and 18th levels have 92 and 163 linear parameters, respectively, and four radial functions for each. Therefore, the number of the radial parameters can be excessive and some of them are not necessary for an accurate fitting. In Figure  6 only the values for MF with the compression of 50% became greater than the ones for MF with the optimal ranks. In general, the optimal ranks and the ranks related to the compression of 50% are close to each other (see Figures  2 (a, b, c) and  3) for the potentials fitted on Mo-Nb-Ta-W, and therefore the results demonstrated in Figures  5 and  6 are close to each other.
In Figures  7 and  8 we see the values of the loss functions and the fitting errors for FLiNaK. From Figure  7 we conclude that the MF, MFRA, and TF MTPs of all levels with the optimal ranks give smaller loss functions and force errors than the base MTPs. However, R-MF demonstrates values similar to the base MTPs and gives greater force fitting error and loss function for the 12th level. Energy fitting errors are statistically similar for all models except for R-MF of the 12th and 16th levels, yielding the worst energy errors. In Figure  8 we observe similar loss functions and energy errors for all MTPs excluding R-MF of the 12th and 16th levels, again demonstrating the worst accuracy among all the models considered. At the same time, force errors are still lower for MF, MFRA, and TF of levels greater than 12 compared to the base MTP even for a two times reduced number of radial parameters.
@@
Figure 7: Loss function and root mean square errors (RMSEs) for energies and forces for potentials trained on FLiNaK with optimal ranks. We provide the results with 68% confidence interval (i.e., 1-$\sigma$ interval).
@@
Figure 8: Loss function and root mean square errors (RMSEs) for energies and forces for potentials trained on FLiNaK with 50% compression. We provide the results with 68% confidence interval (i.e., 1-$\sigma$ interval).
@@
Finally, we fitted ensembles of different types of 20th-level MTPs on the training set including glycine molecular crystals. We used the same level of MTP as in Ref. Rybin, Novikov, and Shapeev (2025a) and we also took the training set from this paper. The results are shown in Table  2. We conclude that the potentials with the optimal ranks give slightly smaller loss functions and fitting errors than the potentials with the compression of 50% of the radial parameters, but the difference is negligible: approximately 0.5 meV/atom for energy, less than 10 meV/Å for forces, and less than 0.1 eV for stresses. We also note that the ensemble of the base MTPs and the ensembles of the compressed potentials with the optimal ranks yield similar loss functions, energy and force errors, and these compressed potentials give slightly smaller stress errors. Thus, as for Mo-Nb-Ta-W and FLiNaK, reducing the number of the radial parameters does not worsen the accuracy of potentials fitted on glycine.
@@
base - 149 $\pm$ 6 5.2 $\pm$ 0.3 109 $\pm$ 2 1.06 $\pm$ 0.03
MF opt 9 137 $\pm$ 6 5.2 $\pm$ 0.2 105 $\pm$ 2 0.95 $\pm$ 0.05
MF 50% 6 159 $\pm$ 15 5.7 $\pm$ 0.4 112 $\pm$ 5 1.03 $\pm$ 0.05
R-MF opt 11 147 $\pm$ 8 5.3 $\pm$ 0.2 109 $\pm$ 3 1.02 $\pm$ 0.03
R-MF 50% 7 172 $\pm$ 17 5.8 $\pm$ 0.4 117 $\pm$ 5 1.07 $\pm$ 0.07
MFRA opt 9 126 $\pm$ 3 4.9 $\pm$ 0.2 100 $\pm$ 1 0.96 $\pm$ 0.04
MFRA 50% 6 144 $\pm$ 4 5.3 $\pm$ 0.2 107 $\pm$ 1 0.99 $\pm$ 0.03
TF opt 4,10,5 133 $\pm$ 6 5.1 $\pm$ 0.2 103 $\pm$ 2 0.93 $\pm$ 0.04
TF 50% 4,8,4 144 $\pm$ 8 5.4 $\pm$ 0.1 107 $\pm$ 3 0.98 $\pm$ 0.04
From the figures and the table provided in this subsection, we make several important conclusions. First, compressed MTPs of the 14th level and higher with optimal ranks improve the accuracy of the base MTPs and potentials with the compression of 50% of the radial parameters do not significantly degrade the accuracy of the base potentials. Second, the accuracy of MTPs with the twice reduced number of the radial parameters is worse, but close to the MTPs with the optimal ranks. Therefore, an MTP with compression of 50% is a reasonable compromise between the accuracy and the number of parameters. Finally, the MFRA potentials recommended themselves as the most accurate MTPs in terms of the loss function values and fitting errors. Thus, the rank augmentation provided in Algorithm  1 is a useful procedure in terms of the accuracy of fitting.
@@
Figure 9: Relative stability of glycine polymorphs calculated using MTP of the 20th level with different compression algorithms.
III.4 Temperature dependence of the density of molten FLiNaK
To examine the predictive power and stability of the compressed MTPs at finite temperature here, we calculated the density of molten LiF-NaF-KF (in an eutectic composition 46.5-11.5-42 mol% usually labeled as FLiNaK). FLiNaK at temperatures of 800-1200 K. To that end, we conducted molecular dynamics simulations (MD) for 200 picoseconds in the NPT ensemble for the system of 1512 atoms. We utilized the LAMMPS package Thompson et al. (2022) for MD simulations. From Tabel 3, we conclude that all potentials, including those with compression of 50% of the radial parameters, give similar densities, which are in good agreement with the experimental result Romatoski and Hu (2017) and previous calculations with the base MTP Rybin et al. (2024).
III.5 Energy ranking of glycine polymorphs
@@
We utilized training datasets obtained in a previous studyRybin, Novikov, and Shapeev (2025a). Our benchmarks with the base and compressed potentials, yielding the smallest values of the loss function, demonstrate that the three most stable known polymorphs of glycine ($\alpha$, $\beta$, and $\gamma$) are correctly identified as the most energetically favorable. In all cases (even in the compression by 50%), the MTP-based rankings show $\alpha$-glycine to have the lowest enthalpy (see Figure 9), with the $\beta$ and $\gamma$ phases being slightly higher in energy. This result is consistent with prior DFT-based studies Zhu et al. (2012), though it contrasts with the experimental relative thermodynamic stability of $\gamma$ > $\alpha$ > $\beta$. This discrepancy is not unexpected and points to the need for more accurate methods of computing intermolecular interaction energies Chalykh et al. (2025).
III.6 Discussion
The main goal of this work was to verify the methods developed for compression of MLIPs on the example of MTP. We demonstrated the robustness of the proposed methods and showed that even 50% compression of the radial parameters of MTP retains the accuracy of the base MTP. To calculate the gradients of the loss function with respect to the radial parameters of the compressed MTPs, we used the gradients of the loss function with respect to the radial parameters of the base MTP. Thus, we did not explicitly implement the compressed MTPs, but it can be explicitly implemented in the program code. The profit from the explicit program realization of the compressed form of MTPs is the computational cost of simulations with these potentials. They have fewer parameters than the base MTPs and, therefore, they require less computational resources.
@@ IV Conclusions
The methodology is verified using Moment Tensor Potential (MTP) model – an example of widely used MLIP. We presented all derivations and benchmarked the methodology on three systems: Mo-Nb-Ta-W, molten FLiNaK, and a glycine molecular crystal. The MTP model was selected for testing due to its wide applicability and well-established benchmark performance. Our results showed that even with compression of up to 50% of the radial parameters of the base (non-compressed) MTP, the fitting errors remain comparable and are sometimes even lower. In principle, as demonstrated, compressing up to 50% of the radial parameters presents a reasonable compromise between accuracy and the number of parameters, indicating that half of the radial MTP parameters are excessive, at least for high-level potentials.
@@
In the case of FLiNaK, the compressed MTPs are robust, i.e., suitable for large-scale MD simulations, and reproduce the melt densities at finite temperatures with accuracy comparable to the base MTP. In contrast to the Mo-Nb-Ta-W system, compressing 50% of the radial parameters was not optimal for the MTPs fitted on FLiNaK; consequently, only the force fitting errors for the MF, TF, and MFRA models were smaller than those of the base MTP.
@@ Acknowledgements.
Supplementary Material: Low-rank matrix and tensor approximations: advancing efficiency of machine-learning interatomic potentials
@@ IV.1 Basics of Riemannian manifolds
For Riemannian manifolds, a tangent space $T_{X}\mathcal{M}$ is defined at each point $X$, representing a local linearization of the neighborhood around that point, see Figure 1 from the main text for illustration. When the manifold $\mathcal{M}$ is unambiguous, we shorten the notation for the tangent space from $T_{X}\mathcal{M}$ to simply $T_{X}$. Formally, the tangent space at $X\in\mathcal{M}_{r}$ is given by Ref. Absil, Mahony, and Sepulchre (2009):
@@
$\operatorname{grad}\leavevmode\nobreak\ f(X)=\operatorname{Proj}_{T_{X}}\nabla f(X),$
@@
For optimization, it is sometimes necessary to transfer vectors from one tangent space to another. This requires a vector transport operator. The illustration of the concept is shown on Figure  10. In this work, we use the operator $\mathcal{T}_{X\to Y}:T_{X}\mathcal{M}_{r}\to T_{Y}\mathcal{M}_{r}$, chosen such that for all $\xi\in T_{X}$, $\mathcal{T}_{X\to Y}\xi=\operatorname{Proj}_{T_{Y}}\xi$.
@@ IV.2 Riemannian BFGS
Now consider a specific case in which we are interested in this work: the manifold $\mathcal{M}_{r}$ of fixed rank $r$ matrices of size $m\times n$. As we mentioned earlier, sometimes we think of matrices as of vectors implying natural isomorphism $\operatorname{vec}(\cdot):\mathbb{R}^{m\times n}\to\mathbb{R}^{mn}$. Let $\tilde{\mathcal{M}}_{r}$ be a manifold of vectorized matrices from $\mathcal{M}_{r}$. For matrices $X,Y$ we denote $x=\operatorname{vec}(X),\leavevmode\nobreak\ y=\operatorname{vec}(Y)$.
@@
where $P_{U_{Y}}^{\perp}=I_{m}-(U_{Y})_{r}(U_{Y})_{r}^{\top},\leavevmode\nobreak\ P_{V_{Y}}^{\perp}=I_{n}-(V_{Y})_{r}(V_{Y})_{r}^{\top}$ and $E$ is a matrix consisting only of 0 and 1 such that Multiplying by this matrix “cuts out” special columns of the initial matrix.
@@
For $\mathcal{M}_{r}$, the explicit form of matrices in $T_{X}\mathcal{M}_{r}$ (see Ref. Steinlechner (2016)) is:
@@
where $X=U_{X}\Sigma_{X}V_{X}^{\top}$ – full SVD and $U_{X,r}=(U_{X})_{r},\leavevmode\nobreak\ V_{X,r}=(V_{X})_{r}$. Show that the set of matrices lying in $T_{X}\mathcal{M}_{r}$ and having a form $U_{X}E_{ij}V_{X}^{\top}$, where $i\leq r$ or $j\leq r$, is orthogonal:
@@
Notice that the cardinality of this set is (n+mr)r(n+m-r)r. It equals dimensionality of $T_{X}\mathcal{M}_{r}$, therefore, $\{U_{X}E_{ij}V_{X}^{\top}\mid i\leq r\leavevmode\nobreak\ \text{or}\leavevmode\nobreak\ j\leq r\}$ is an orthogonal matrix basis in $T_{X}\mathcal{M}_{r}$ and $\{\operatorname{vec}(U_{X}E_{ij}V_{X}^{\top})\mid i\leq r\leavevmode\nobreak\ \text{or}\leavevmode\nobreak\ j\leq r\}$ is an orthogonal vector basis in $T_{x}\tilde{\mathcal{M}}_{r}$. To obtain $B_{x}$ we have to stack these vectors:
@@
Recall that we denote $P_{U_{X}}^{\perp}=(I_{m}-U_{X,r}U_{X,r}^{\top}),\leavevmode\nobreak\ P_{V_{X}}^{\perp}=(I_{n}-V_{X,r}V_{X,r}^{\top})$.
@@
Figure 12: Histogram with the ranks and the number of parameters (linear and radial) for potentials of the 12th and 16th levels fitted on the F-Li-Na-K training set.

### v2

@@ III.3 Comparison of the compressed MTPs and ETN for Mo-Nb-Ta-W
Figure 5: Root mean square errors (RMSEs) for energies and forces calculated with ETN and the compressed MTPs on the Mo-Nb-Ta-W validation set. We provide the results with 68% confidence interval (i.e., 1-$\sigma$ interval).
III.4 Loss functions and training errors
@@
In this subsection, we analyze the values of loss functions and fitting errors calculated with ensembles of different MTPs with the optimal ranks and with the compression of approximately 50% (i.e., two times reduction) of the number of the radial parameters in the base MTP. In addition, on the example of the FLiNaK system we demonstrate that the tensor of the ACE radial parameters can be compressed with matrix factorization.
@@
The results of fitting of potentials on Mo-Nb-Ta-W are shown in Figures 6 and 7. In Figure 6 we observe the decrease in loss functions, energy and force fitting errors with the increase of the level of all MTPs. We also see that all the compressed MTPs (MF, MFRA, R-MF, and TF) of the 16th and 18th levels give higher accuracy, i.e., smaller loss functions, energy and force fitting errors than the base MTPs. However, for the MTPs of the 12th level, we conclude that the compressed MTPs give either worse (R-MF) or slightly lower (MF, TF, and MFRA) results compared to the base MTP. This effect is related to the number of linear parameters and the complexity of MTP. There are 29 linear parameters and three radial functions in the MTP of the 12th level, and thus, all the radial parameters presented in the base MTP can be critical for accurate fitting of the compressed MTPs. Furthermore, the MTPs of the 6th and 10th level have two and three radial functions, respectively, and they only have 5 and 16 linear parameters, respectively. As demonstrated in the supplementary material, compression of these potentials gave slightly greater values of the loss functions for the MF MTP at any rank and, thus, there is no sense in compression of the MTPs of these levels from the point of view of accuracy. High-level MTPs of the 16th and 18th levels have 92 and 163 linear parameters, respectively, and four radial functions for each. Therefore, the number of the radial parameters can be excessive and some of them are not necessary for an accurate fitting. In Figure 7 only the values for MF with the compression of 50% became greater than the ones for MF with the optimal ranks. In general, the optimal ranks and the ranks related to the compression of 50% are close to each other (see Figures 2 (a, b, c) and 3) for the potentials fitted on Mo-Nb-Ta-W, and therefore the results demonstrated in Figures 6 and 7 are close to each other.
@@
Figure 6: Loss function and root mean square errors (RMSEs) for energies, forces, stresses for different MTPs trained on Mo-Nb-Ta-W with the optimal ranks. The calculations are conducted on the validation set. We provide the results with 68% confidence interval (i.e., 1-$\sigma$ interval).
@@
Figure 7: Loss function and root mean square errors (RMSEs) for energies, forces, stresses for different potentials trained on Mo-Nb-Ta-W with 50% compression. The calculations are conducted on the validation set. We provide the results with 68% confidence interval (i.e., 1-$\sigma$ interval).
It should be noted that the MFRA MTP and the MF MTP require the same CPU time for their fitting. From Table 2 we see that MF and MFRA require from 540 to 555 CPU minutes to be fitted and the fitting time is particularly independent of rank. The MF MTP and the MFRA MTP require the same CPU time, as we use the same number of BFGS iterations to optimize each of them, and most of the BFGS steps for the MFRA MTP are conducted at $r=r_{\rm max}$ (see Algorithm 1 for details), which almost corresponds to fitting of the MF MTP at the same, fixed rank. Thus, we do not sacrifice CPU time for accurate training of the MFRA MTP.
Table 2: Total CPU time (in minutes) required for fitting of the MF MTP and the MFRA MTP of the 14th level with various ranks on the Mo-Nb-Ta-W system. The results are given for the ensembles of five MF and MFRA with 68% confidence interval (i.e., 1-$\sigma$ interval).
$r=4$ $r=6$ $r=8$ $r=10$
MF $542\pm 6$ $543\pm 9$ $544\pm 6$ $545\pm 16$
MFRA ($\Delta r=2$) $45\pm 5 55$$2\pm 2 55$$2\pm 12 55$$55\pm 18$
In Figures 8 and 9 we see the values of the loss functions and the fitting errors for FLiNaK. From Figure 8 we conclude that the MF, MFRA, and TF MTPs of all levels with the optimal ranks give smaller loss functions and force errors than the base MTPs. However, R-MF demonstrates values similar to the base MTPs and gives greater force fitting error and loss function for the 12th level. Energy fitting errors are statistically similar for all models except for R-MF of the 12th and 16th levels, yielding the worst energy errors. In Figure 9 we observe similar loss functions and energy errors for all MTPs excluding R-MF of the 12th and 16th levels, again demonstrating the worst accuracy among all the models considered. At the same time, force errors are still lower for MF, MFRA, and TF of levels greater than 12 compared to the base MTP even for a two times reduced number of radial parameters.
Figure 8: Loss function and root mean square errors (RMSEs) for energies, forces, stresses for potentials trained on FLiNaK with optimal ranks. We provide the results with 68% confidence interval (i.e., 1-$\sigma$ interval).
Figure 9: Loss function and root mean square errors (RMSEs) for energies, forces, stresses for potentials trained on FLiNaK with 50% compression. We provide the results with 68% confidence interval (i.e., 1-$\sigma$ interval).
In Figure 10 the energy and force fitting errors are shown for the ensembles of the base and the compressed ACE fitted on the FLiNaK training set. It could be seen that we do not lose in accuracy when we compress the radial part of ACE and reduce the number of the radial parameters 1.5 times using the factorization of the reshaped matrix of the ACE radial parameters. On the example of the FLiNaK system we see that the proposed methodology for compression of MLIPs is universal and can be applied to various potentials.
Figure 10: Root mean square errors (RMSEs) for energies and forces for the base and compressed ACE potentials trained on FLiNaK. We provide the results with 68% confidence interval (i.e., 1-$\sigma$ interval).
@@
Finally, we fitted ensembles of different types of 20th-level MTPs on the training set including glycine molecular crystals. We used the same level of MTP as in Ref. Rybin, Novikov, and Shapeev (2025) and we also took the training set from this paper. The results are shown in Table 3. We conclude that the potentials with the optimal ranks give slightly smaller loss functions and fitting errors than the potentials with the compression of 50% of the radial parameters, but the difference is negligible: approximately 0.5 meV/atom for energy, less than 10 meV/Å for forces, and less than 0.1 eV for stresses. We also note that the ensemble of the base MTPs and the ensembles of the compressed potentials with the optimal ranks yield similar loss functions, energy and force errors, and these compressed potentials give slightly smaller stress errors. Thus, as for Mo-Nb-Ta-W and FLiNaK, reducing the number of the radial parameters does not worsen the accuracy of potentials fitted on glycine.
@@
base - 148 $\pm$ 14 5.2 $\pm$ 0.4 109 $\pm$ 5 1.05 $\pm$ 0.08
MF opt 9 134 $\pm$ 6 5.1 $\pm$ 0.2 104 $\pm$ 2 0.95 $\pm$ 0.04
MF 50% 6 159 $\pm$ 13 5.7 $\pm$ 0.4 112 $\pm$ 4 1.06 $\pm$ 0.07
R-MF opt 11 149 $\pm$ 8 5.2 $\pm$ 0.2 109 $\pm$ 3 1.02 $\pm$ 0.03
R-MF 50% 7 172 $\pm$ 16 5.8 $\pm$ 0.5 117 $\pm$ 5 1.07 $\pm$ 0.06
MFRA opt 9 130 $\pm$ 4 5.0 $\pm$ 0.2 102 $\pm$ 2 0.98 $\pm$ 0.04
MFRA 50% 6 146 $\pm$ 8 5.4 $\pm$ 0.2 108 $\pm$ 3 1.01 $\pm$ 0.04
TF opt 4,10,5 129 $\pm$ 6 4.9 $\pm$ 0.2 101 $\pm$ 2 0.92 $\pm$ 0.05
TF 50% 4,8,4 143 $\pm$ 7 5.4 $\pm$ 0.2 107 $\pm$ 3 0.99 $\pm$ 0.03
From the figures and the table provided in this subsection, we make several important conclusions. First, compressed MTPs of the 14th level and higher with optimal ranks improve the accuracy of the base MTPs and potentials with the compression of 50% of the radial parameters do not significantly degrade the accuracy of the base potentials. Second, the accuracy of MTPs with the twice reduced number of the radial parameters is worse, but close to the MTPs with the optimal ranks. Therefore, an MTP with compression of 50% is a reasonable compromise between the accuracy and the number of parameters. Finally, the MFRA potentials recommended themselves as the most accurate MTPs in terms of the loss function values and fitting errors. Thus, the rank augmentation provided in Algorithm 1 is a useful procedure in terms of the accuracy of fitting.
@@
Figure 11: Relative stability of glycine polymorphs calculated using MTP of the 20th level with different compression algorithms and DFT. Data obtained in the scope of DFT is taken from our previous calculations Rybin, Novikov, and Shapeev (2025). Averaging is done over 10 separately trained potentials for each model.
III.5 Temperature dependence of the density of molten FLiNaK
To examine the predictive power and stability of the compressed MTPs at finite temperature here, we calculated the density of molten LiF-NaF-KF (in an eutectic composition 46.5-11.5-42 mol% usually labeled as FLiNaK). FLiNaK at temperatures of 800-1200 K. To that end, we conducted molecular dynamics simulations (MD) for 200 picoseconds in the NPT ensemble for the system of 1512 atoms. We utilized the LAMMPS package Thompson et al. (2022) for MD simulations. From Tabel 4, we conclude that all potentials, including those with compression of 50% of the radial parameters, give similar densities, which are in good agreement with the experimental result Romatoski and Hu (2017) and previous calculations with the base MTP Rybin et al. (2024).
III.6 Energy ranking of glycine polymorphs
@@
We utilized training datasets obtained in a previous studyRybin, Novikov, and Shapeev (2025). Our benchmarks with the base and compressed MTPs demonstrate that each MTP (even with the 50% compression) identified $\alpha$-glycine to have the lowest energy (see Figure 11), whereas the $\beta$ and $\gamma$ phases have slightly higher energies. This result is consistent with prior DFT-based studies Zhu et al. (2012), though it contrasts with the experimental relative thermodynamic stability of $\gamma$ > $\alpha$ > $\beta$. This discrepancy is not unexpected and points to the need for more accurate methods of computing intermolecular interaction energies Chalykh et al. (2025).
III.7 Discussion
The main goal of this work was to verify the methods developed for compression of MLIPs, particularly MTP. We demonstrated the robustness of the proposed methods and showed that even 50% compression of the radial parameters of MTP retains the accuracy of the base MTP. To calculate the gradients of the loss function with respect to the radial parameters of the compressed MTPs, we used the gradients of the loss function with respect to the radial parameters of the base MTP. Thus, we did not explicitly implement the compressed MTPs, but it can be explicitly implemented in the program code. The profit from the explicit program realization of the compressed form of MTPs is the computational cost of simulations with these potentials. They have fewer parameters than the base MTPs and, therefore, they may require less computational resources for creation of a training set for their fitting.
@@ IV Conclusions
The methodology is mainly verified using Moment Tensor Potential (MTP) model – an example of widely used MLIP. We presented all derivations and benchmarked the methodology on three systems: Mo-Nb-Ta-W, molten FLiNaK, and a glycine molecular crystal. The MTP model was selected for testing due to its wide applicability and well-established benchmark performance. Our results showed that even with compression of up to 50% of the radial parameters of the base (non-compressed) MTP, the fitting errors remain comparable and are sometimes even lower. In principle, as demonstrated, compressing up to 50% of the radial parameters presents a reasonable compromise between accuracy and the number of parameters, indicating that half of the radial MTP parameters are excessive, at least for high-level potentials.
@@
In the case of FLiNaK, the compressed MTPs are robust, i.e., suitable for large-scale MD simulations, and reproduce the melt densities at finite temperatures with accuracy comparable to the base MTP. In contrast to the Mo-Nb-Ta-W system, compressing 50% of the radial parameters was not optimal for the MTPs fitted on FLiNaK; consequently, only the force fitting errors for the MF, TF, and MFRA models were smaller than those of the base MTP. In addition to the compressed MTPs, we developed the compressed Atomic Cluster Expansion (ACE) and tested it on the FLiNaK system. It was demonstrated that the compressed ACE requires 1.5 times less radial parameters to be fitted as accurately as the base (original) ACE. Thus, the proposed methodology for reduction of a number of MLIPs parameters is universal and can be applied to different MLIPs.
@@ Acknowledgements.
Supplementary Material: Low-rank matrix and tensor approximations for compression of machine-learning interatomic potentials
@@ IV.1 Basics of Riemannian manifolds
For Riemannian manifolds, a tangent space $T_{X}\mathcal{M}$ is defined at each point $X$, representing a local linearization of the neighborhood around that point, see Figure 1 from the main text for illustration. When the manifold $\mathcal{M}$ is unambiguous, we shorten the notation for the tangent space from $T_{X}\mathcal{M}$ to simply $T_{X}$. Formally, the tangent space at $X\in\mathcal{M}_{r}$ is given by Absil, Mahony, and Sepulchre (2009):
@@
$\operatorname{grad}\penalty 10000\ f(X)=\operatorname{Proj}_{T_{X}}\nabla f(X),$
@@
For optimization, it is sometimes necessary to transfer vectors from one tangent space to another. This requires a vector transport operator. The illustration of the concept is shown on Figure 12. In this work, we use the operator $\mathcal{T}_{X\to Y}:T_{X}\mathcal{M}_{r}\to T_{Y}\mathcal{M}_{r}$, chosen such that for all $\xi\in T_{X}$, $\mathcal{T}_{X\to Y}\xi=\operatorname{Proj}_{T_{Y}}\xi$.
@@ IV.2 Riemannian BFGS
Now consider a specific case in which we are interested in this work: the manifold $\mathcal{M}_{r}$ of fixed rank $r$ matrices of size $m\times n$. As we mentioned earlier, sometimes we think of matrices as of vectors implying natural isomorphism $\operatorname{vec}(\cdot):\mathbb{R}^{m\times n}\to\mathbb{R}^{mn}$. Let $\tilde{\mathcal{M}}_{r}$ be a manifold of vectorized matrices from $\mathcal{M}_{r}$. For matrices $X,Y$ we denote $x=\operatorname{vec}(X),\penalty 10000\ y=\operatorname{vec}(Y)$.
@@
where $P_{U_{Y}}^{\perp}=I_{m}-(U_{Y})_{r}(U_{Y})_{r}^{\top},\penalty 10000\ P_{V_{Y}}^{\perp}=I_{n}-(V_{Y})_{r}(V_{Y})_{r}^{\top}$ and $E$ is a matrix consisting only of 0 and 1 such that Multiplying by this matrix “cuts out” special columns of the initial matrix.
@@
For $\mathcal{M}_{r}$, the explicit form of matrices in $T_{X}\mathcal{M}_{r}$Steinlechner (2016) is:
@@
where $X=U_{X}\Sigma_{X}V_{X}^{\top}$ – full SVD and $U_{X,r}=(U_{X})_{r},\penalty 10000\ V_{X,r}=(V_{X})_{r}$. Show that the set of matrices lying in $T_{X}\mathcal{M}_{r}$ and having a form $U_{X}E_{ij}V_{X}^{\top}$, where $i\leq r$ or $j\leq r$, is orthogonal:
@@
Notice that the cardinality of this set is (n+mr)r(n+m-r)r. It equals dimensionality of $T_{X}\mathcal{M}_{r}$, therefore, $\{U_{X}E_{ij}V_{X}^{\top}\mid i\leq r\penalty 10000\ \text{or}\penalty 10000\ j\leq r\}$ is an orthogonal matrix basis in $T_{X}\mathcal{M}_{r}$ and $\{\operatorname{vec}(U_{X}E_{ij}V_{X}^{\top})\mid i\leq r\penalty 10000\ \text{or}\penalty 10000\ j\leq r\}$ is an orthogonal vector basis in $T_{x}\tilde{\mathcal{M}}_{r}$. To obtain $B_{x}$ we have to stack these vectors:
@@
Recall that we denote $P_{U_{X}}^{\perp}=(I_{m}-U_{X,r}U_{X,r}^{\top}),\penalty 10000\ P_{V_{X}}^{\perp}=(I_{n}-V_{X,r}V_{X,r}^{\top})$.
@@
Gradients of the loss function with respect to parameters of tensor train
Gradients of the loss function $L(\Xi,G^{(1)},G^{(2)},G^{(3)},G^{(4)})$ with respect to all tensor train cores $G^{(k)},k=1,\ldots,4$:
@@
Figure 14: Histogram with the ranks and the number of parameters (linear and radial) for potentials of the 12th and 16th levels fitted on the FLiNaK training set.
MF MTP of the 6th and 10th level
From Figure 15 we see that the values of the loss function of the MF MTPs are greater than the loss function of the base MTP. Thus, it has no sense to compress MTPs of these levels from the point of view of accuracy.
Figure 15: Dependence of the loss function calculated on the Mo-Nb-Ta-W validation set on the rank of the MF MTP of (a) the 6th level and of (b) the 10th level. Error bars indicate 1-$\sigma$ confidence intervals.
Recommendations for choice of MFRA hyperparameters
Setup.
We study MFRA training on the Mo-Nb-Ta-W training set for potentials of the 14th level by varying three hyperparameters: (1) the augmentation interval $s$, (2) the initial rank $r_{\min}$, and (3) the rank increment $\Delta r$. Unless otherwise noted, we use a total budget of 2000 BFGS steps with $\Delta r=2,s=80,r_{\min}=2$ and provide the value of the loss function on the validation set.
Findings.
  1. 1.
    Step interval $s$. The loss function is broadly insensitive to $s$ across a wide range, with a shallow minimum around $s\in[60,100]$ for the 2000-step budget (Figure 16a). Too small $s$ (e.g., $\leq 20$) induces MF-like behavior because the ranks grow before the optimizer adapts the low-rank subspace. In contrast, a very large $s$ ineffectively uses the computing budget at higher ranks.
  2. 2.
    Initial rank $r_{\min}$. The initial rank $r_{\min}=1$ consistently underperforms higher ranks (Figure 16b). Choosing $r_{\min}\in\{2,3,4\}$ yields smaller and similar values of the loss functions. An excessively large $r_{\min}$ can skip a helpful low-rank warm-up and mildly worsen the final value of the loss function.
  3. 3.
    Rank increment $\Delta r$. $\Delta r=1$ performs poorly due to too many effective restarts; all the other choices presented for this parameter ($2\leq\Delta r\leq 4$) demonstrate similar behavior and give reasonable values of the loss function (Figure 16c).
Recommendations.
  1. 1.
    Use $s$ as 3–5% of the total iteration budget. For the $2000$ iterations, this corresponds to $s\approx 60$100100.
  2. 2.
    Do not take $r_{\min}=1$; prefer $r_{\min}\geq 2$ for a stable warm-up without wasting computational time.
  3. 3.
    Do not use $\Delta r=1$; we suggest using $\Delta r\simeq r_{\min}$.
(a) Augmentation interval $s$.
(b) Initial rank $r_{\min}$.
(c) Rank increment $\Delta r$.
Figure 16: Dependence of the loss function on hyperparameters for 14th-level MFRA MTPs on the Mo–Nb–Ta–W validation set: (a) the augmentation interval $s$, (b) the initial rank $r_{\min}$, and (c) the rank increment $\Delta r$. Error bars indicate 1-$\sigma$ confidence intervals.